# Segre forms of singular metrics on vector bundles and Lelong numbers

### v1

@@ 4.4. Multiplicities of quasi-cycles
Let $\xi$ be a tuple of holomorphic functions that defines the maximal ideal at $x\in Y$. If $\mu\in{\mathcal{QZ}}(Y)$ and $u=\log|\xi|^{2}$, then $M^{u}\wedge\mu$ is in ${\mathcal{QZ}}(Y)$ and has support at $x$. Thus it has the form $\alpha[x]$ for some integer $\alpha$, cf. Example 4.5. By Proposition 4.16 (ii) and (4.21) it follows that $\alpha$ is independent of the choice of tuple $\xi$ defining the maximal ideal. If $\mu$ has bidegree $(k,k)$, for degree reasons, we thus have that $M^{u}\wedge\mu=M^{u}_{n-k}\wedge\mu$.
@@
Let $\mu\in{\mathcal{QZ}}(Y)$ be a $(k,k)$-current and $x\in Y$. We define the multiplicity of $\mu$ at $x$, $\operatorname{mult}_{x}\mu$, as
@@
If $\mu\in{\mathcal{QZ}}_{0}(Y)$ then $M^{u}\wedge\mu\in{\mathcal{QZ}}_{0}(Y)$ by Proposition 4.16 (i), and hence $\operatorname{mult}_{x}\mu=0$, cf. (4.21). That is,
@@
Remark 4.20.
If $\mu$ is a positive closed current, it is clear from (4.23), that $\operatorname{mult}_{x}\mu$ is the Lelong number of $\mu$ at $x$. ∎
@@
Assume that $\mu\in{\mathcal{QZ}}(Y)$. Then any multiplicities $\operatorname{mult}_{x}\mu$ is an integer. If $\mu$ has bidegree $(k,k)$, then there is a unique decomposition
@@
where $\mu_{F}$ is the Lelong current of an analytic cycle on $Y$ and the set where $\operatorname{mult}_{x}\mu_{M}$ does not vanish is contained in an analytic set of codimension $\geq k+1$.
@@
Indeed, it follows from Example 4.5 that $\operatorname{mult}_{x}\mu$ is an integer. Since we know that $\operatorname{mult}_{x}\mu$ is integer valued, (4.24) follows directly from the Siu decomposition in case $\mu$ is a positive current. As in the case with $\mu\in{\mathcal{GZ}}(Y)$, if $\mu=\tau_{*}\gamma$, and $\tau=i\circ\tau^{\prime}$, where $\tau^{\prime}\colon W\to Z$ is surjective, then $\operatorname{mult}_{x}\tau^{\prime}_{*}\gamma=\operatorname{mult}_{i(x)}\tau_{*}\gamma$. Moreover, [1, Lemma 7.3] holds, with the same proof, for $\mu=\tau_{*}\gamma$ if $\gamma$ is as in (4.1). In fact, the crucial point that $M^{\tau^{*}\xi}\wedge\gamma=0$ if $M^{\tau^{*}\xi}=M^{\tau^{*}\xi}\wedge 1=0$. To see this we can assume that $\xi$ is a tuple of functions (i.e., a section of a vector bundle with trivial metric) in a neighborhood of $x$. Then $M^{\tau^{*}\xi}$ is a positive current. If $b_{\ell}$ is locally of the form $u_{\ell}+g_{\ell}$, where $u_{\ell}$ is psh, and $g_{\ell}$ is smooth, we may take a sequence $b_{\ell}^{j}=u_{\ell}^{j}+g_{\ell}$, where each $u_{\ell}^{j}$ is psh, and $u_{\ell}^{j}$ decreases to $u_{\ell}$. Then, by [13, Proposition III.4.9], $M^{\tau^{*}\xi}\wedge\gamma_{j}\to M^{\tau^{*}\xi}\wedge\gamma$. If $M^{\tau^{*}\xi}=0$, thus $M^{\tau^{*}\xi}\wedge\gamma_{j}=0$ and hence $M^{\tau^{*}\xi}\wedge\gamma=0$. Now Theorem 4.21 follows precisely as [1, Theorem 7.1]. ∎
@@
Assume that $\mu\in{\mathcal{QZ}}(Y)$ has bidegree $(p,p)$ and has support on an analytic subspace $V\subset Y$ of pure (co)dimension. Since $\mu$ is a normal current, it follows that $\mu$ is an analytic cycle a priori by real coefficients if ${\rm codim\,}V=p$, and $\mu=0$ if ${\rm codim\,}V>p$. Since the multiplicities are integral we see that it is actually a cycle with integer coefficients, i.e., in ${\mathcal{Z}}(Y)$. If $\mu$ is in ${\mathcal{QZ}}_{0}(Y)$, then all its multiplicities vanish, and hence it is 00. Thus ${\mathcal{Z}}(Y)$ is indeed a submodule of ${\mathcal{QB}}(Y)$. ∎
@@ 5. Regularization of qas metrics on a line bundle
Let $L\to Y$ be a holomorphic vector bundle, $e^{\psi_{0}}$ a smooth reference metric on $L$, and let $\omega=s_{1}(L,e^{\psi_{0}})$. Note that a singular metric $e^{\psi}$ on $L$ is a qas if and only if the global function $\psi-\psi_{0}$ is qas. Indeed, given a local trivialization $L\cong X\times{\mathbb{C}}_{\alpha}$,
$\psi-\psi_{0}=\log|\alpha|^{2}_{e^{\psi}}-\log|\alpha|^{2}_{e^{\psi_{0}}}=\log|1|^{2}_{e^{\psi}}-\log|1|^{2}_{e^{\psi_{0}}},$
so if the metric $e^{\psi}$ is qas, then the right-hand side is qas by Lemma 2.1. The other direction follows by writing
$\log|\alpha|^{2}_{e^{\psi}}=(\psi-\psi_{0})+\log|\alpha|^{2}_{e^{\psi_{0}}},$
and using that $\log|\alpha|^{2}_{e^{\psi_{0}}}$ is qas.
We define the endomorphism
(5.1) $[dd^{c}\psi]_{\omega}\wedge\mu=[dd^{c}(\psi-\psi_{0})]\wedge\mu+\omega\wedge\mu$
on ${\mathcal{QZ}}(Y)$. It follows from Proposition 4.8 that (5.1) indeed maps ${\mathcal{QZ}}(Y)$ to ${\mathcal{QZ}}(Y)$.
Remark 5.1.
Assume that $L$ is trivial and $\psi_{0}=0$. Then the notation is consistent if $[dd^{c}\psi]=[dd^{c}\psi]_{0}$. ∎
Remark 5.2.
If $Z=Y$, i.e., $\psi\equiv-\infty$, it is natural to let $[dd^{c}\psi]_{\omega}\wedge\mu=\omega\wedge\mu$ for all $\mu$. ∎
@@
(i) If $e^{\psi}$ is a qas metric on $L\to Y$, then $[dd^{c}\psi]^{k}_{\omega}\colon{\mathcal{QZ}}_{0}(Y)\to{\mathcal{QZ}}_{0}(Y)$.
(ii) If $e^{\psi^{\prime}}$ is another qas metric and $\psi^{\prime}=\psi+{\mathcal{O}}(1)$ then $[dd^{c}\psi^{\prime}]^{k}_{\omega}-[dd^{c}\psi]^{k}_{\omega}\colon{\mathcal{QZ}}(Y)\to{\mathcal{QZ}}_{0}(Y)$.
@@
(5.2) $[dd^{c}\psi]_{\omega}\wedge\mu=\tau_{*}\big{(}[dd^{c}\tau^{*}\psi]_{\tau^{*}\omega}\wedge\mu^{\prime}\big{)}.$
@@
Although $[dd^{c}\psi]_{\omega}$ seems to be the fundamental endomorphism, it is sometimes useful to express it in terms of $[dd^{c}\psi]$ which is well-defined although $\psi$ is only a metric, not a global function. Let $Z$ be the unbounded locus of $e^{\psi}$, and let $\omega=dd^{c}\psi_{0}$. If $\mu={\bf 1}_{Y\setminus Z}\mu$, then (5.1) is
$[dd^{c}(\psi-\psi_{0})]\wedge\mu+\omega\wedge\mu=:[dd^{c}\psi]\wedge\mu,$
and if $\mu={\bf 1}_{Z}\mu$, then it is just $\omega\wedge\mu$. Therefore,
(5.3) $[dd^{c}\psi]_{\omega}\wedge\mu=[dd^{c}\psi]\wedge{\bf 1}_{Y\setminus Z}\mu+\omega\wedge{\bf 1}_{Z}\mu,$
which coincides with how the corresponding operator is defined in [18]. We also have endomorphisms
@@
(5.4) $[dd^{c}\psi]^{k}_{\omega}\wedge\mu=\langle dd^{c}\psi\rangle^{k}\wedge\mu+\sum_{\ell=0}^{k}\omega^{k-\ell}\wedge M^{\psi}_{\ell}\wedge\mu,\quad k=0,1,2,\ldots.$
@@
Since $\psi-\psi_{0}$ is a global function, the left hand side in (5.5) is well-defined. The smoothness assumption on $\chi_{j}$ is not necessary, but for simplicity we restrict to this case, since we will only apply the theorem to $\chi_{\epsilon}(t)=\log(e^{t}+{\epsilon})$.
The theorem is known when $\mu$ is $1$, however expressed with the right hand side in (5.4). In [7] it is proved for a more general class of (q)psh functions. Here we will rely on the presentation in [10] and show that the proof given there works even when acting on a quasi-cycle $\mu$.
@@
Since the statement is local, we can restrict to an open subset where we have a global frame element, and hence that $\psi$ and $\psi_{0}$ are fixed functions. For simplicity we let $Y$ denote this open set.
We first consider the case when $\psi_{0}=0$. If $\mu$ has support on $Z$, then both sides of (5.5) are equal to $\omega^{k}\mu$. Thus we can assume that $\mu=\tau^{*}\gamma$, where $\tau\colon W\to Y$, $\gamma$ is a product as in (4.1) and $\tau^{-1}Z$ has positive codimension in $W$. After possibly another modification we may assume that $\tau^{*}\psi=\log|\sigma|^{2}+a$, where $\sigma$ is a section of a Hermitian line bundle ${\mathcal{L}}\to W$ and $a$ is qpsh and locally bounded. We claim that
@@ 6. Segre forms and endomorphisms of qas bundles
Assume that $(E,h)$ is a qas bundle over $X$. If $e^{\psi}$ denotes the induced metric on $L_{E}$, then $\psi=\log|\alpha|^{2}_{h}$ is qas, i.e., for any local frame $a$ of $L_{E}$, i.e., the well-defined function $\log|\alpha a|^{2}$ is qas. Note that on the set where $\alpha_{j}\neq 0$ we may take $a=1/\alpha_{j}$.
@@
If $Z^{\prime}\subset{\mathbb{P}}(E)$ denotes the set where $\psi$ is not locally bounded, then we have outside of $Z^{\prime}$ that $s_{1}(L_{E},e^{\psi})=dd^{c}\psi$. Recall, cf. (5.1) and (5.3), that
(6.1) $[dd^{c}\psi]_{\omega}\wedge\mu=dd^{c}(\psi{\bf 1}_{X\setminus Z^{\prime}}\mu)+\omega\wedge{\bf 1}_{Z^{\prime}}\mu$
defines an endomorphism on ${\mathcal{QZ}}({\mathbb{P}}(E))$. In view of Lemma 4.12 and Proposition 4.16 (iii) the following definition makes sense.
@@
Given a qas bundle $(E,h)$ over $X$ we have the Segre endomorphisms ${\mathcal{QZ}}(X)\to{\mathcal{QZ}}(X)$ and ${\mathcal{QZ}}_{0}(X)\to{\mathcal{QZ}}_{0}(X)$ of degree $k$, defined by the formula
(6.2) ${\bf s}_{k}(E,h,h_{0})\wedge\mu=p_{*}([dd^{c}\psi]_{\omega}^{k+r-1}\wedge p^{*}\mu),$
and ${\bf s}(E,h,h_{0}):=\sum_{k}{\bf s}_{k}(E,h,h_{0})$.
If $h$ is smooth, $[dd^{c}\psi]_{\omega}$ is just multiplication with the smooth form $dd^{c}\psi$, so ${\bf s}_{k}(E,h,h_{0})$ is just multiplication by the smooth form $s_{k}(E,h)$, see (4.16).
We define the Segre form as $s(E,h,h_{0})={\bf s}(E,h,h_{0})1$ on $X$, so that
(6.3) $s(E,h,h_{0}):=p_{*}\sum_{k=0}^{\infty}p_{*}[dd^{c}\psi]^{k}_{\omega}.$
Note that if $L$ is a line bundle, then $p_{L}:{\mathbb{P}}(L)\to X$ is the identity. Hence,
@@
The definition of $s(E,h,h_{0})$ and ${\bf s}(E,h,h_{0})$ here coincides with the definition in [18] except that in those articles is used the convention that ${\mathbb{P}}(E)_{x}$ denotes the hyperplanes in $E_{x}$ so one has to make some necessary adaptions. Indeed, in [18], the main results are formulated for singular metrics $h$ which are positively curved, in contrast to the negatively curved metrics we consider here. This difference is however essentially artificial. For a singular metric $h$, there is an induced dual metric $h^{*}$ on the dual bundle $E^{*}$, see e.g., [17, Lemma 3.1]. By definition, a singular metric $h$ on $E$ is Griffiths positive if $h^{*}$ is Griffiths negative. The definition of Segre form in [18] is such that if $h$ is a Griffiths positive singular metric on $E$, then it holds just as in the smooth case that
@@
where now the left-hand side is the Segre form defined in [18], and the right-hand side is the Segre form defined in this article. ∎
Below, we will sometimes drop $h$ and $h_{0}$ from the notation of the Segre forms and operators, and consider $E$ to be a qas bundle, i.e., equipped with a qas metric $h$, and furthermore a smooth reference metric $h_{0}$.
@@
(6.6) ${\bf s}(E,h,h_{0})\wedge\mu=\sum_{\ell=0}^{\infty}p_{*}(\langle dd^{c}\psi\rangle^{\ell}\wedge p^{*}\mu)+p_{*}\big{(}s(L_{E},e^{\psi_{0}})\wedge M^{\psi}\wedge p^{*}\mu\big{)}.$
We may decompose ${\bf s}_{k}(E,h,h_{0})\wedge\mu$ as
${\bf s}_{k}^{\prime}(E,h,h_{0})\wedge\mu+M^{E,h,h_{0}}\wedge\mu:={\bf 1}_{X\setminus Z}{\bf s}_{k}(E,h,h_{0})\wedge\mu+{\bf 1}_{Z}{\bf s}_{k}(E,h,h_{0})\wedge\mu.$
By (6.6), it follows that
$s_{k}^{\prime}(E,h,h_{0})\wedge\mu=p_{*}(\langle dd^{c}\psi\rangle^{\ell}\wedge p^{*}\mu)\text{ and }M^{E,h,h_{0}}_{k}\wedge\mu=s(L_{E},e^{\psi_{0}})\wedge M^{\psi}\wedge p^{*}\mu.$
In particular, it follows that $s^{\prime}(E,h,h_{0})$ is actually independent of $h_{0}$, and may hence be denoted simply by $s^{\prime}(E,h)$.

### v2

@@ 4.4. Multiplicities of quasi-cycles
Let $\xi$ be a tuple of holomorphic functions that defines the maximal ideal at $x\in Y$. If $\mu\in{\mathcal{QZ}}(Y)$ and $u=\log|\xi|^{2}$, then $M^{u}\wedge\mu$ is in ${\mathcal{QZ}}(Y)$ and has support at $x$, so it is $\alpha[x]$ for some integer $\alpha$, cf. Example 4.5. By Proposition 4.16 (ii) and Proposition 4.15, $\alpha$ is independent of the choice of tuple $\xi$ defining the maximal ideal. If $\mu$ has bidegree $(k,k)$, for degree reasons, we have that $M^{u}\wedge\mu=M^{u}_{n-k}\wedge\mu$.
@@
Let $\mu\in{\mathcal{QZ}}(Y)$ be a $(k,k)$-current and $x\in Y$. We define the multiplicity of $\mu$ at $x$, $\operatorname{mult}_{x}\mu$, as the integer
@@
If $\mu$ is a positive closed current, (4.23) means that $\operatorname{mult}_{x}\mu$ is the Lelong number of $\mu$ at $x$.
If $\mu\in{\mathcal{QZ}}_{0}(Y)$, then $M^{u}\wedge\mu\in{\mathcal{QZ}}_{0}(Y)$ by Proposition 4.16 (i), and hence $\operatorname{mult}_{x}\mu=0$, cf. Proposition 4.15. That is,
@@
Assume that $\mu\in{\mathcal{QZ}}(Y)$. If $\mu$ has bidegree $(k,k)$, then there is a unique decomposition
@@
where $\mu_{F}$ is the Lelong current of an analytic cycle on $Y$ of codimension $k$, and the set where $\operatorname{mult}_{x}\mu_{M}$ does not vanish is contained in an analytic set of codimension $\geq k+1$.
@@
Since we know that $\operatorname{mult}_{x}\mu$ is integer valued, (4.24) follows directly from the Siu decomposition in case $\mu$ is a positive current. As in the case with $\mu\in{\mathcal{GZ}}(Y)$, if $\mu=\tau_{*}\gamma$, and $\tau=i\circ\tau^{\prime}$, where $\tau^{\prime}\colon W\to Z$ is surjective, then $\operatorname{mult}_{x}\tau^{\prime}_{*}\gamma=\operatorname{mult}_{i(x)}\tau_{*}\gamma$. Moreover, [1, Lemma 7.3] holds, with the same proof, for $\mu=\tau_{*}\gamma$ if $\gamma$ is as in (4.1). In fact, the crucial point is that $M^{\tau^{*}\xi}\wedge\gamma=0$ if $M^{\tau^{*}\xi}=M^{\tau^{*}\xi}\wedge 1=0$. To see this we can assume that $\xi$ is a tuple of functions (i.e., a section of a vector bundle with trivial metric) in a neighborhood of $x$. Then $M^{\tau^{*}\xi}$ is a positive current. If $b_{\ell}$ is locally of the form $u_{\ell}+g_{\ell}$, where $u_{\ell}$ is psh, and $g_{\ell}$ is smooth, we may take a sequence $b_{\ell}^{j}=u_{\ell}^{j}+g_{\ell}$, where each $u_{\ell}^{j}$ is psh, and $u_{\ell}^{j}$ decreases to $u_{\ell}$. Then, by [13, Proposition III.4.9], $M^{\tau^{*}\xi}\wedge\gamma_{j}\to M^{\tau^{*}\xi}\wedge\gamma$. If $M^{\tau^{*}\xi}=0$, thus $M^{\tau^{*}\xi}\wedge\gamma_{j}=0$ and hence $M^{\tau^{*}\xi}\wedge\gamma=0$. Now Theorem 4.20 follows precisely as [1, Theorem 7.1]. ∎
@@
Assume that $\mu\in{\mathcal{QZ}}(Y)$ has bidegree $(k,k)$ and support on an analytic subspace $i\colon V\to Y$ of pure codimension $q$. In view of Proposition 4.4, $\mu=i_{*}\tau$ for a $\tau\in{\mathcal{QZ}}(V)$ of bidegree (kq,kq)(k-q,k-q). Thus $\mu=0$ if $k<q$. If $k=q$, then $\tau$ is a locally constant function which is integer-valued since its multiplicities are integers. It follows that $\mu$ is a cycle, i.e., in ${\mathcal{Z}}(Y)$. If $\mu$ is in ${\mathcal{QZ}}_{0}(Y)$, then it vanishes by Proposition 4.15. Thus ${\mathcal{Z}}(Y)$ is a submodule of ${\mathcal{QB}}(Y)$. ∎
@@ 5. Regularization of qas metrics on a line bundle
Let $L\to Y$ be a holomorphic vector bundle, let $e^{\psi}$ be a qas metric on $L$, and let $e^{\psi_{0}}$ be a smooth reference metric. Then, cf. Example 3.2, the global function $\psi-\psi_{0}$ is qas on $Y$. In view of Proposition 4.8 the following definition makes sense.
Definition 5.1.
Consider a qas metric $e^{\psi}$ on $L\to X$, a smooth reference metric $e^{\psi_{0}}$, and $\omega=dd^{c}\psi_{0}$. We introduce the endomorphism
(5.1) $dd^{c}\psi]_{\omega}\wedge\mu=[dd^{c}(\psi-\psi_{0})]\wedge\mu+\omega\wedge\mu$
on ${\mathcal{QZ}}(Y)$.
Recall that if $\psi\equiv$-$\infty$, then $dd^{c}(\psi-\psi_{0})]\wedge\mu=0$, so that $dd^{c}\psi]_{\omega}=\omega\wedge\mu$.
@@
(i) If $e^{\psi}$ is a qas metric on $L\to Y$, then $dd^{c}\psi]^{k}_{\omega}\colon{\mathcal{QZ}}_{0}(Y)\to{\mathcal{QZ}}_{0}(Y)$.
(ii) If $e^{\psi^{\prime}}$ is another qas metric and $\psi^{\prime}=\psi+{\mathcal{O}}(1)$ then $dd^{c}\psi^{\prime}]^{k}_{\omega}-[dd^{c}\psi]^{k}_{\omega}\colon{\mathcal{QZ}}(Y)\to{\mathcal{QZ}}_{0}(Y)$.
@@
(5.2) $dd^{c}\psi]_{\omega}\wedge\mu=\tau_{*}\big([dd^{c}\tau^{*}\psi]_{\tau^{*}\omega}\wedge\mu^{\prime}\big).$
@@
It is sometimes useful to express $dd^{c}\psi]_{\omega}$ in terms of $dd^{c}\psi]$ which is well-defined even when $\psi$ is a metric. In fact, (4.9) is local and not affected if $u$ is replaced by $u+h$, where $h$ is pluriharmonic. Let $Z$ be the degeneracy locus of $e^{\psi}$. If $\mu={\bf 1}_{Y\setminus Z}\mu$, then (5.1) is $dd^{c}(\psi-\psi_{0})]\wedge\mu+\omega\wedge\mu=[dd^{c}\psi]\wedge\mu,$ and if $\mu={\bf 1}_{Z}\mu$, then it is just $\omega\wedge\mu$. Therefore,
(5.3) $dd^{c}\psi]_{\omega}\wedge\mu=[dd^{c}\psi]\wedge{\bf 1}_{Y\setminus Z}\mu+\omega\wedge{\bf 1}_{Z}\mu;$
this is precisely the definition of $dd^{c}\psi]_{\omega}$ in [18]. Besides $dd^{c}\psi]^{k}$ we have endomorphisms
@@
(5.4) $dd^{c}\psi]^{k}_{\omega}\wedge\mu=\langle dd^{c}\psi\rangle^{k}\wedge\mu+\sum_{\ell=0}^{k}\omega^{k-\ell}\wedge M^{\psi}_{\ell}\wedge\mu,\quad k=0,1,2,\ldots.$
@@
Since $\psi-\psi_{0}$ is a global function $\psi_{j}$ are well-defined. The smoothness assumption on $\chi_{j}$ is not necessary, but for simplicity we restrict to this case, since we will only apply the theorem to $\chi_{\epsilon}(t)=\log(e^{t}+{\epsilon})$.
This theorem is known when $\mu$ is $1$ and with $dd^{c}\psi]^{k}_{\omega}$ expressed as in (5.4). In [7] it is proved for a more general class of (q)psh functions. Here we will rely on the presentation in [10] and show that the proof given there works even when acting on a quasi-cycle $\mu$.
@@
First assume that $\psi\equiv$-$\infty$. Then $\psi_{j}=\chi_{j}(-\infty)+\psi_{0}$ so that $(dd^{c}\psi_{j})^{k}\wedge\mu=\omega^{k}\wedge\mu=[dd^{c}\psi]^{k}_{\omega}\wedge\mu$. Let us thus assume that $\psi\not\equiv$-$\infty$. Since the statement is local, we can restrict to an open subset where we have a global frame element, and hence that $\psi$ and $\psi_{0}$ are fixed functions. For simplicity we let $Y$ denote this open set.
We begin with the case when $\psi_{0}=0$. If $\mu$ has support on $Z$, then both sides of (5.5) are equal to $\omega^{k}\wedge\mu$. Thus we can assume that $\mu=\tau^{*}\gamma$, where $\tau\colon W\to Y$, $\gamma$ is a product as in (4.1) and $\tau^{-1}Z$ has positive codimension in $W$. After possibly another modification we may assume that $\tau^{*}\psi=\log|\sigma|^{2}+a$, where $\sigma$ is a section of a Hermitian line bundle ${\mathcal{L}}\to W$ and $a$ is qpsh and locally bounded. We claim that
@@ 6. Segre forms and endomorphisms of qas bundles
Assume that $(E,h)$ is a qas bundle over $X$. If $e^{\psi}$ denotes the induced metric on $L_{E}$, then $\psi=\log|\alpha|^{2}_{h}$ is qas. Thus for any local frame $a$ of $L_{E}$ the function $\log|\alpha a|^{2}_{h}$ is qas on ${\mathbb{P}}(E)$. On the set where $\alpha_{j}\neq 0$ we may take $a=1/\alpha_{j}$.
@@
If $Z^{\prime}\subset{\mathbb{P}}(E)$ denotes the set where $\psi$ is not locally bounded, then outside of $Z^{\prime}$ we have that $s_{1}(L_{E},e^{\psi})=dd^{c}\psi$. Recall, cf. (4.9) and (5.3), that
(6.1) $dd^{c}\psi]_{\omega}\wedge\mu=dd^{c}(\psi{\bf 1}_{X\setminus Z^{\prime}}\mu)+\omega\wedge{\bf 1}_{Z^{\prime}}\mu$
defines an endomorphism on ${\mathcal{QZ}}({\mathbb{P}}(E))$. In view of Lemma 4.12 and Proposition 5.2 (i) the following definition makes sense.
@@
Given a qas bundle $(E,h)$ over $X$, and a smooth metric $h_{0}$, we have the Segre endomorphisms ${\mathcal{QZ}}(X)\to{\mathcal{QZ}}(X)$ and ${\mathcal{QZ}}_{0}(X)\to{\mathcal{QZ}}_{0}(X)$ of bidegree $(k,k)$,
(6.2) ${\bf s}_{k}(E,h,h_{0})\wedge\mu=p_{*}([dd^{c}\psi]_{\omega}^{k+r-1}\wedge p^{*}\mu).$
We let
${\bf s}(E,h,h_{0}):=\sum_{k}{\bf s}_{k}(E,h,h_{0}).$
We define the Segre form $s(E,h,h_{0}):={\bf s}(E,h,h_{0})1$ on $X$, i.e.,
(6.3) $s(E,h,h_{0}):=p_{*}\sum_{k=0}^{\infty}[dd^{c}\psi]^{k}_{\omega}.$
Note that if $L$ is a line bundle, then $p_{L}\colon{\mathbb{P}}(L)\to X$ is the identity. Hence,
@@
The definition of $s(E,h,h_{0})$ here coincides with the definition in [18]. However, in that article is used the convention that ${\mathbb{P}}(E)_{x}$ denotes the hyperplanes in $E_{x}$ so one has to make some necessary adaptions. Moreover, the definition is made on the associated (Griffiths positive) dual bundle $(E^{*},h^{*})$, see, e.g., [17, Lemma 3.1]. As in the smooth case one have the relations
@@
which connect the definitions in this article and in [18]. ∎
@@
(6.6) ${\bf s}(E,h,h_{0})\wedge\mu=p_{*}\big(\sum_{\ell=0}^{\infty}\langle dd^{c}\psi\rangle^{\ell}\wedge p^{*}\mu+s(L_{E},e^{\psi_{0}})\wedge M^{\psi}\wedge p^{*}\mu\big).$
Since the support of $M^{\psi}\wedge p^{*}\mu$ is contained in $p^{-1}Z$,
(6.7) ${\bf 1}_{X\setminus Z}{\bf s}(E,h,h_{0})\wedge\mu=p_{*}\big({\bf 1}_{{\mathbb{P}}(E)\setminus p^{-1}Z}\sum_{\ell=0}^{\infty}\langle dd^{c}\psi\rangle^{\ell}\wedge p^{*}\mu\big).$
It is independent of choice of reference metric $h_{0}$ and we denote it by ${\bf s}^{\prime}(E,h)\wedge\mu$. Letting
$M^{E,h,h_{0}}\wedge\mu:={\bf 1}_{Z}{\bf s}_{k}(E,h,h_{0})\wedge\mu$
we have the decomposition
(6.8) ${\bf s}(E,h,h_{0})\wedge\mu={\bf s}_{k}^{\prime}(E,h)\wedge\mu+M^{E,h,h_{0}}\wedge\mu.$
Note that by (4.6), both terms on the right-hand side belong to ${\mathcal{QZ}}(X)$.
@@
6.1. The case when $h$ is qas and non-degenerate
We need the following technical result.